OpenStudy (anonymous):

plot the graph of y=3x^3 - 2x^2 +x - 10 for the values of x from -10 to 10 using matlab command and apply Newton Raphson method to find the positive root of the equation y=3x^3 - 2x^2 +x - 10 using matlab

OpenStudy (anonymous):

then i guess you need matlab!

OpenStudy (experimentx):

for plotting x = -10:0.1:10; plot(x, 3x^3 - 2x^2 +x - 10)

OpenStudy (experimentx):

can you explain is short sentence how newtonw raphson method works??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its like lagrangian interpolation

OpenStudy (zarkon):

I wouldn't say that

OpenStudy (experimentx):

looks like this is getting interesting

OpenStudy (zarkon):

Is there a newtions method command in matlab?...or are you supposed to program one?

OpenStudy (anonymous):

i dont need the matlab help, i need manual help...using newton raphson method

OpenStudy (zarkon):

I assume you know the iterative formula?

OpenStudy (anonymous):

For Newton Raphson Method First find f(x) then using the equation x_n+1 = x_n - f(x_n) / f'(x_n) find the required formulae

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

do you have your initial guess \(x_0\)?

OpenStudy (anonymous):

no i havent attempted yet

OpenStudy (zarkon):

ok...where exactly are you stuck?

OpenStudy (experimentx):

so how do we use it to find the roots?

OpenStudy (anonymous):

pick a number you think will be close to a solution to \[3x^3 - 2x^2 +x - 10=0\]

OpenStudy (anonymous):

that is your first guess, say \[x_0\] second guess will be \[x_0-\frac{f(x_0)}{f'(x_0)}\]

OpenStudy (experimentx):

i guess i understand it not ... hahahah

OpenStudy (experimentx):

*now

OpenStudy (anonymous):

if you pick \[x_0=\frac{5}{3}\] you are done in one step since \[f(\frac{5}{3})=0\]

OpenStudy (anonymous):

if you pick say \[x_0=2\] then your second guess will be \[2-\frac{f(2)}{f'(2)}\]

OpenStudy (anonymous):

\[f(2)=8\] \[f'(x)=6x^2-4x+1\] \[f'(2)=17\] \[2-\frac{8}{17}=\frac{26}{17}\] is your second guess

OpenStudy (experimentx):

is there some kind of error here function xn = function_newt(x) xn = x - (3*x^3 - 2*x^2 +x - 10)/(9*x^2 - 4*x +1);

OpenStudy (zarkon):

satellite73 just has a small typo in his derivative.

OpenStudy (experimentx):

x = 3; for i = 0:10 x = function_newt(x); end disp(x)

OpenStudy (experimentx):

function xn = function_newt(x) xn = x - (3*x^3 - 2*x^2 +x - 10)/(9*x^2 - 4*x +1);